OpenStudy (anonymous):

1. x – 3 > –9 2. –2x + 1 ≤ –11 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3) 6. Explain, in complete sentences, when you would use an open circle or a closed circle, and when you would shade to the right or left, to graph an inequality on the number line.

OpenStudy (anonymous):

7. Describe a real life scenario where inequalities are used and state the inequality.

OpenStudy (anonymous):

1. x - 3 > - 9 x > -9 + 3 x > - 6 2. -2x + 1 <= - 11 -2x <= - 11 - 1 -2x <= - 12 x => -12/-2 x => 6 3. 10 < -3x + 1 10 - 1 < - 3x 9 < - 3x 9/-3 > x -3 > x 4. 2(x + 5) > 8x - 8 2x + 10 > 8x - 8 2x - 8x > -8 - 10 -6x > - 18 x < -18/-6 x < 3 5. -2(x - 3) >= 5 - (x + 3) -2x + 6 >= 5 -x - 3 -2x + x >= 2 - 6 -x >= - 4 x <= -4/-1 x <= 4 6. To graph an inequality on a number line : An open circle is used if the inequality sign is greater then or less then. If there is an equal sign, then you do not use an open circle. A closed circle is used when an equal sign is present. It can be used if there is a greater then or equal to , or less then or equal to. For instance, if you had x < 8 or x > 8, you would use an open circle. If your inequality looked like this, x <=4 or x >=4, you would then use a closed circle. As far as the shading goes....x < 4...this is saying that x is less then 4. If it is less then, the shading goes to the left. If the inequality stated x > 4...x is greater then 4, then the shading would go to the right. example : x < 5 -- open circle, shading to the left. x >= 5 -- closed circle, shading to the right x > 10 -- open circle, shading to the right x <= 3 -- closed circle, shading to the left. 7. You are going to the store to buy some fruit. You have $5. You want to purchase apples and bananas. a = apples and b = bananas a + b <=5 Basically, all this is saying is that the apples plus the bananas has to be less then or equal to $5 in order for you to be able to purchase it without going over.